d2y/dx2 - dy/dx= e-x , y(0) =0, y'(0) =1'

1
Expert's answer
2021-06-21T17:29:49-0400
"y''-y'=e^{-x}"

Write the related homogeneous or complementary equation:


"y''-y'=0"


The general solution of a nonhomogeneous equation is the sum of the general solution "y_h(x)" of the related homogeneous equation and a particular solution "y_p(x)" of the nonhomogeneous equation:


"y(x)=y_h(x)+y_p(x)"

Consider a homogeneous equation 


"y''-y'=0"

Write the characteristic (auxiliary) equation:


"r^2-r=0"

"r(r-1)=0"

"r_1=0, r_2=1"

The general solution of the homogeneous equation is


"y_h(x)=C_1+C_2e^{x}"



Find the particular solution of the nonhomogeneous differential equation


"y_p(x)=Ae^{-x}"

"y_p'=-Ae^{-x}"

"y_p''=Ae^{-x}"

Substitute


"Ae^{-x}-(-Ae^{-x})=e^{-x}"

"A=\dfrac{1}{2}"

"y_p(x)=\dfrac{1}{2}e^{-x}"

The general solution of a second order homogeneous differential equation is


"y(x)=C_1+C_2e^{x}+\dfrac{1}{2}e^{-x}"


Initial conditions: "y(0)=0, y'(0)=1"



"y(0)=C_1+C_2e^{0}+\dfrac{1}{2}e^{-0}=0"

"y'(x)=C_2e^{x}-\dfrac{1}{2}e^{-x}"

"y'(0)=C_2e^{0}-\dfrac{1}{2}e^{-0}=1"

"C_2=\dfrac{3}{2}"

"C_1=-2"

The solution of the given differential equation is


"y(x)=-2+\dfrac{3}{2}e^{x}+\dfrac{1}{2}e^{-x}"